# Anti-Phishing Training Does Not Work: A Large-Scale Empirical Assessment of Multi-Modal Training Grounded in the NIST Phish Scale

### v1

@@ Abstract
Social engineering attacks using email, commonly known as phishing, are a critical cybersecurity threat. Phishing attacks often lead to operational incidents and data breaches. As a result, many organizations allocate a substantial portion of their cybersecurity budgets to phishing awareness training, driven in part by compliance requirements. However, the effectiveness of this training remains in dispute. Empirical evidence of training (in)effectiveness is essential for evidence-based cybersecurity investment and policy development. Despite recent measurement studies, two critical gaps remain in the literature: (1) we lack a validated measure of phishing lure difficulty, and (2) there are few comparisons of different types of training in real-world business settings.
To fill these gaps, we conducted a large-scale study ($N=12{,}511$) of phishing effectiveness at a US-based financial technology (“fintech”) firm. Our two-factor design compared the effect of treatments (lecture-based, interactive, and control groups) on subjects’ susceptibility to phishing lures of varying complexity (using the NIST Phish Scale). The NIST Phish Scale successfully predicted behavior (click rates: 7.0% easy to 15.0% hard emails, p $<$ 0.001), but training showed no significant main effects on clicks (p = 0.450) or reporting (p = 0.417). Effect sizes remained below 0.01, indicating little practical value in any of the phishing trainings we deployed. Our results add to the growing evidence that phishing training is ineffective, reinforcing the importance of phishing defense-in-depth and the merit of changes to processes and technology to reduce reliance on humans, as well as rebuking the training costs necessitated by regulatory requirements.
@@ 1 Introduction
Phishing attacks critically impact organizations through operational incidents, data breaches, and financial losses. They constitute the primary attack vector to breach organizations as highlighted by industry studies [1] and the US Federal Bureau of Investigation (FBI) [2, 1]. To prevent this class of attacks, organizations employ technical controls (e.g., spam filters, attachment analyzers) complemented by cybersecurity awareness training [3, 4].
Although these anti-phishing trainings are costly, their effectiveness is in question. Organizational security awareness programs often focus on compliance metrics rather than meaningful behavioral change [5]. Meanwhile, recent large-scale evidence suggests that current approaches have limited impact [4]. To better evaluate the effectiveness of anti-phishing training, two questions arise. First, are some kinds of training more effective than others? Second, how does their effectiveness vary with the difficulty of the phishing lure?
In this paper, we report the results of a measurement study to address these questions. We situate the study in an operational environment [6], set at a US-based international fintech firm with 12,511 employees across multiple business units. We compared three experimental conditions: a control group (no training), traditional lecture-based training, and lectures augmented with interactive phishing exercises. Subjects in each condition received lures of varying complexity, which we assigned difficulty levels using the NIST Phish Scale. This study design enables rigorous evaluation of training effectiveness across standardized difficulty levels on a large population, improving on previous works that had either been small-scale or lacking in systematic difficulty assessment [7, 8].
Our findings were unfavorable for the conventional anti-phishing trainings available from cybersecurity training vendors. Neither mode of training produced statistically significant improvements in phish click reduction (average: 10.4%) or phish reporting behavior (average: 9.6%). However, the NIST Phish Scale did predict user behavior, with phishing lure difficulty having a highly significant effect on click-through rates (F(2, 12086) = 41.415, p $<$ 0.001, $\eta^{2}$ = 0.007). Thus our study shows that, while the NIST Phish Scale provides a means of measuring the difficulty of different lures, the trainings we applied did not help subjects perform better on those lures.
@@
    Large-Scale Validation of the NIST Phish Scale. We provide the most comprehensive empirical validation of the NIST Phish Scale to date, demonstrating its predictive utility across 12,511 participants in an operational environment, showing strong correlation between difficulty ratings and user susceptibility.
@@
    Realistic Assessment of Training Effectiveness. Our results challenge optimistic assumptions about the effectiveness of phishing training. We found no statistically significant impact of training on either click rates or reporting behavior. Although these findings align with recent large-scale studies from the healthcare [4] and banking sectors [9], our study addresses explicit calls from the research community for large-scale ecologically valid studies combining standardized assessments of difficulty inside of an operational context [10, 6].
Significance: Our findings have important implications for researchers, organizations, and regulators. For the research community, we provide empirical assessment of training effectiveness using rigorous methodology and standardized difficulty measures. For organizations, our results suggest the need for realistic expectations about training outcomes and highlight the importance of complementing awareness programs with robust technical controls rather than relying primarily on human-centered defenses. Our validation of the NIST Phish Scale’s predictive utility enables organizations to use this tool for benchmarking the difficulty of their phishing simulations. For regulators, our work contributes to the growing body of evidence that organizational cybersecurity strategy should emphasize technical defenses [1] rather than placing the burden on humans — requiring anti-phishing training is costly and has minimal effect.
@@ 2 Background and Related Work
In a phishing attack, attackers send messages (phishing lures) to targets whom they wish to persuade to act against their own interests. Attackers seek results such as sensitive information, account credentials, or the ability to control the victim’s machine [1, 11]. Modern phishing campaigns target individuals and organizations across multiple attack vectors, applying conventional email phishing as well as SMS-based attacks (“smishing”) and voice-based deception (“vishing”) [12, 13]. Recent attacks have begun to incorporate Generative AI tools to produce tailored and realistic messages. The severity, persistence, and increasing sophistication of these attacks has established phishing detection and prevention as a critical focus within both academic cybersecurity research and industry practice [14, 6].
In this section, we summarize current defenses against phishing (Appendix 2.1), discuss metrics for evaluating phishing effectiveness (Appendix 2.2), and review prior empirical studies (Appendix 2.3) and knowledge gaps (Appendix 2.4).
2.1 Phishing Defenses
Phishing poses substantial risks to organizations, who therefore erect multiple layers of defenses against these attacks. We distinguish automated defenses from human-focused ones.
@@ 2.1.1 Automated Defenses
Organizations employ a range of technological defenses to mitigate phishing risks. These include spam filters, attachment analyzers, and URL checkers [15, 14, 1]. More recently, machine learning approaches have been applied to detect phishing content. Long Short-Term Memory (LSTM) models and optimized Random Forest algorithms [16] have demonstrated high accuracy in detecting phishing attempts. As with everything in cybersecurity, phishing is a cat-and-mouse game. Adversaries have begun to leverage artificial intelligence to create more convincing lures. Large language models (LLMs) can generate personalized and context-aware phishing emails that are difficult for traditional filters to detect [3]. Techniques like typosquatting, domain spoofing, and watering hole attacks exacerbate this issue [1, 14].
2.1.2 Human-Focused Defenses: Training and Awareness
Defending against phishing is understood not as solely a technical problem, but rather as a socio-technical one [6, 13]. Spear phishing and emerging phishing attacks present sophisticated threats that overcome automated defenses [17, 18], so human awareness of phishing remains necessary to successful defense [19, 20].
@@ 2.1.2 Human-Focused Defenses: Training and Awareness
There are many approaches to training humans to be resilient to phishing attacks. Methods range from lecture-style training involving passive videos and readings [21] to simulation feedback providing immediate feedback after interaction with phishing simulations [22, 23]. Interactive training uses gamification or scenario-based learning to engage participants [24, 25], while just-in-time training alerts users in real time during risky behavior [26].
@@
Anti-phishing trainings are frequently included within the cybersecurity awareness trainings ubiquitous in organizations. Legal mandates drive the widespread adoption of such training programs, with HIPAA requiring security awareness training for healthcare organizations [27], GDPR encouraging staff training on cybersecurity practices [28], and PCI DSS mandating training for those handling cardholder data [29]. Industry standards and frameworks further reinforce these requirements, with ISO 27001 calling for human resource security controls including awareness programs [30] and the NIST Cybersecurity Framework advising phishing simulations under the “Identify” function [31].
2.2 Measuring Effectiveness of Phishing Training
The effectiveness of phishing training can be assessed based on how resilient a target becomes to phishing lures. Appendix 2.2.1 describes how lures can be categorized by their complexity, and then Appendix 2.2.2 gives the standard metrics for target resilience.
2.2.1 Measuring Training Lure Difficulty
TABLE I: NIST Phish Scale for Detection Difficulty [32, 33].
@@ 2.2.1 Measuring Training Lure Difficulty
TABLE II: Summary of Prior Experiments on Phishing and Training Interventions. Studies are organized by research context (Lab vs. Real-World). Context: “Lab” denotes research simulations, where subjects are not in their work roles. “Real-World” denotes studies where participants were employees acting in their job capacity. Complexity measure: Whether the study controlled for the difficulty of the phishing tasks.  indicates the use of an open control such as NIST Phish Scale, and  indicates no control. Hypotheses: How these works informed the hypotheses in the present study (cf. Table III).
[8] (2024) Lab University students 96 H3
[34] (2024) Lab University students 117  (NIST Phish Scale) H4
[7] (2022) Real-world Employees 119 H2
[22] (2017) Real-world University students 150 H2, H3
[23] (2014) Real-world Employees 1,359 H2, H4
[9] (2024) Real-world Banking employees “Thousands” H2
[35] (2023) Real-world Finance employees 5,000 H2
[4] (2025) Real-world Healthcare employees 19,500+ H2
Ours Real-world Fintech org. 12,511  (NIST Phish Scale)
To assess phishing lure difficulty, the NIST Phish Scale categorizes them along two dimensions as depicted in Table I. Phishing Cues represent observable errors or inconsistencies in the email that might alert users to its fraudulent nature, such as spelling mistakes, suspicious URLs, or formatting irregularities [34]. Premise Alignment measures the relevance and familiarity of the email’s subject matter to the recipient’s typical email environment and organizational context [36]. A high alignment, messages that closely resemble emails users receive daily, combined with few obvious cues yields the most deceptive and challenging-to-detect phishing emails [17]. The scale allows organizations to benchmark lure complexity and contextualize simulation outcomes, though it requires sufficiently large sample sizes for reliable use and consistent application [37].
2.2.2 Training Outcome Metrics
The effectiveness of anti-phishing training is commonly measured in three ways:
  • Open Rate captures the percentage of emails opened, though this measurement can be deceptive since different email clients and security settings can affect whether an email is recorded as opened [36] we therefore largely ignore this measure.
  • Click-Through Rate (CTR) measures the percentage of recipients who click on malicious links, serving as a primary indicator of phishing susceptibility [22, 23].
  • Reporting Rate tracks phishing incidents reported to security teams, representing users’ proactive security behaviors and ability to recognize suspicious content [18, 38].
2.3 Empirical Studies of Anti-Phishing Training
For reviews of knowledge about susceptibility to phishing and about anti-phishing training, we refer the reader to recent works by Marshall et al. [21] and Franz et al. [10]. This section focuses on the aspect of the literature to which this study contributes. We discuss the settings of prior work (Appendix 2.3.1), the assessments of different trainings (Appendix 2.3.2), and measurement and evaluation methods (Appendix 2.3.3).
Figure 1: Illustration of our methodology. Our two-factor design randomly assigns different trainings to different subjects, and then randomly sends phishing lures of varying complexity to the subjects. We compared between-subject performance using lure click rate and lure report rate in order to test four hypotheses derived from prior work.
2.3.1 Scale and Context of Prior Work
Many phishing training studies have been conducted in controlled laboratory settings, where researchers benefit from a high degree of experimental control. Studies of this sort have provided numerous insights, including that active learning and immediate feedback can improve short-term detection rates, and have explored how individual differences affect training outcomes [7, 21]. However, as noted by Franz et al. [10], these controlled settings may lack ecological validity. In laboratory environments, participant samples are typically small and consist of homogeneous groups such as academics and students. Due to this, and the artificial nature of lab tasks, these studies may not accurately reflect the complexities and distractions of real-world environments. This raises questions about the generalizability of laboratory findings to operational contexts, where additional factors can alter training effectiveness.
In response, a growing number of studies have begun examining phishing training in real-world organizational environments. Doing et al. [9] analyzed phishing reporting behaviors across thousands of bank employees, while Ho et al. [4] conducted a comprehensive evaluation with over 19,500 healthcare employees across 10 simulated phishing campaigns, reporting minimal effectiveness of both annual cybersecurity awareness training and embedded anti-phishing training. To date, however, these larger studies have lacked a standardized method for assessing the difficulty of the phishing lures in question [4, 9], or have focused primarily on observational analysis rather than controlled training interventions with standardized difficulty measures [9].
Our study contributes to the evidence on real-world human phishing resilience, with data from approximately 12,000 subjects in a fintech enterprise, while maintaining experimental control and measuring lure difficulty with the NIST Phish Scale framework. This approach allows us to examine whether the modest training effects we observe align with Ho et al.’s findings, while providing additional insights into how training effectiveness varies by phishing sophistication.
2.3.2 Training Intervention Approaches
Franz et al. [10] provide a comprehensive taxonomy identifying five categories of phishing interventions: training, awareness, education, notifications, and design interventions. Our study contributes empirical data within their training category, specifically comparing lecture-based delivery against interactive modalities. While Schöni et al. [8] demonstrated effectiveness of personalized training approaches, and our work examines scalable training modalities that can be deployed across demographically diverse cohorts without individualization, recent large-scale studies suggest that the practical benefits of such training interventions may be more limited than previously assumed [4]. Our study builds on this foundation by applying the NIST Phish Scale to measure lure difficulty in line with current guidance, operationalizing the scale at unprecedented scale, and measuring outcomes (clicking, reporting) post-intervention, allowing for practical comparison across conditions and lure complexity.
2.3.3 Measurement and Evaluation Methods
A critical gap in prior work has been the lack of standardized difficulty assessment for phishing simulations. Barrientos et al. [39] conducted early validation work on the NIST Phish Scale, while Dawkins and Jacobs [33] provided the official implementation guidelines. However, no studies have operationalized the scale at enterprise level nor examined how difficulty interacts with training effectiveness. We applied the scale across varied training modalities and measuring multiple outcome measures (clicking, reporting) post-intervention.
2.4 Research Gaps
Table II summarizes the preceding analysis as two gaps. First, prior work has primarily focused on individual-level factors influencing phishing susceptibility. The interaction between training modality and phishing difficulty — a key question for optimizing organizational defenses — has received little attention. Second, despite widespread adoption of phishing training, few studies rigorously assess outcomes against standardized difficulty measures or operate at sufficient scale for robust inference.

### v2

# Anti-Phishing Training (Still) Does Not Work: A Large-Scale Reproduction of Phishing Training Inefficacy Grounded in the NIST Phish Scale

@@ Abstract
Social engineering attacks delivered via email, commonly known as phishing, represent a persistent cybersecurity threat leading to significant organizational incidents and data breaches. Although many organizations train employees on phishing, often mandated by compliance requirements, the real-world effectiveness of this training remains debated.
To contribute to evidence-based cybersecurity policy, we conducted a large-scale reproduction study (N=12,511N=12{,}511) at a US-based financial technology firm. Our experimental design refined prior work by comparing training modalities in operational environments, validating NIST’s standardized phishing difficulty measurement, and introducing novel organizational-level temporal resilience metrics. Echoing prior work, training interventions showed no significant main effects on click rates (pp=0.450) or reporting rates (pp=0.417), with negligible effect sizes. However, we found that the NIST Phish Scale predicted user behavior, with click rates increasing from 7.0% for easy lures to 15.0% for hard lures. Our organizational-level resilience result was mixed: 36-55% of campaigns achieved “innoculation” patterns where reports preceded clicks, but training did not significantly improve organizational-level temporal protection. In summary, our results confirm the ineffectiveness of current phishing training approaches while offering a refined study design for future work.
@@ 1 Introduction
Phishing attacks harm organizations through operational incidents, data breaches, and financial losses [4]. Industry studies and the US FBI agree that phishing is the primary attack vector to breach organizations [42]. As an example, in 2022–2023 campaigns attributed to the Scattered Spider cluster (aka 0ktapus/Muddled Libra) compromised 9,931 accounts in over 130 organizations [43, 23]. This campaign cost MGM Resorts alone an estimated \$100 million [79]. To prevent this class of attacks, organizations employ technical controls (e.g., spam filters, attachment analyzers) complemented by cybersecurity awareness training [58, 48].
Despite widespread organizational investment in anti-phishing training, mounting empirical evidence (and employee sentiment [78]) questions its effectiveness. Large-scale studies across multiple sectors demonstrate minimal training impact [48, 53, 54]. Most pertinently, Lain et al. showed that phishing training did not improve security outcomes across 19,000+ employees over multiple years [53, 54]. These findings align with theoretical work by Volkamer et al., who demonstrate fundamental limitations in human processing of security indicators [96].
This growing body of evidence suggests that phishing training is ineffective, but we believe that further reproduction of the result is needed to inform cybersecurity policy and defense strategies [67, 89]. [^1] Our reproduction effort differs from an exact replication, which would involve repeating identical procedures. Instead we focus on testing the generalizability of the phenomenon of training ineffectiveness, varying the organizational context and the training approach. We target two specific knowledge gaps. First, existing studies lack standardized measures of phishing difficulty, so we do not know whether training effectiveness varies by attack sophistication. Second, most studies examine single training modalities rather than comparing different approaches systematically, so we do not know if some training designs are more effective than others.
To address these limitations, we conducted a large-scale reproduction study (N=12,511N=12{,}511) at a US-based financial technology firm. Our study design has three distinctive elements: (1) mandatory pre-phishing training using two distinct training modalities (lecture-based vs. interactive); (2) systematic difficulty assessment using the NIST Phish Scale [86, 7]; and (3) novel organizational temporal resilience measurement through the Organizational Inoculation Index, which measures the effectiveness of phishing training from an organizational perspective rather than an individual one. This design enables rigorous evaluation of training effectiveness across standardized difficulty levels while providing new insights into organizational cybersecurity resilience patterns that complement individual-focused assessments.
Our reproduction study confirms the ineffectiveness of anti-phishing training across multiple dimensions, but adds nuance to prior results. Consistent with prior work, neither training modality produced statistically significant improvements in phish click reduction (pp=0.450) or phish reporting behavior (pp=0.417), with negligible effect sizes indicating minimal practical value for organizational security investment. However, the NIST Phish Scale successfully predicted user behavior (F(2, 12086) = 41.415, p < 0.001, $\eta^{2}$ = 0.007), providing the first large-scale validation of this standardized difficulty measure. [^2] While the effect size is small in statistical terms due to low base rates of phishing engagement, the practical implications are significant: click rates doubled from 7.0% to 15.0% as lure difficulty increased. Beyond that, our Organizational Inoculation Index revealed substantial organizational resilience that operates independently of individual training effectiveness. Across all experimental conditions, 36.4% to 54.5% of template-group combinations demonstrated “inoculation” patterns where threat reports (from a security-conscious employee) preceded threat interaction by another employee.
@@
    Reproduction of Training Ineffectiveness Across Multiple Modalities. Our results reproduce the core finding of training ineffectiveness while demonstrating this phenomenon’s generalizability across training designs.
@@
    First Large-Scale Validation of the NIST Phish Scale. We evaluate the NIST Phish Scale across 12,511 participants. The Phish Scale is a viable measure for assessing phishing difficulty, with predictive utility.
  3. 3.
    Introduction of the Organizational Inoculation Index. We introduce and validate a novel metric for organizational temporal security. We report substantial organizational resilience (36-55% inoculation rates), which operates independent of individual training effectiveness and provides new insights into collective cybersecurity behaviors.
Significance: Cybersecurity practice and policy must be evidence-based. For regulators, our work contributes to growing evidence that compliance-driven training requirements may not achieve intended security outcomes. For organizations, our results suggest realistic (low) expectations about training outcomes and highlight the importance of technical controls rather than human-centered defenses. However, we also demonstrated that collective security behaviors create measurable protective effects, emphasizing the value of feedback loops between user reports and real-time threat response. For researchers, we validate NIST’s standardized difficulty measure and confirm the reproducibility of training ineffectiveness findings across organizational contexts.
@@ 2 Background and Related Work
Figure 1: Phishing defense architecture, mapping the email pathway from authentication through post-delivery response. Technical controls create multiple interception points prior to human interaction, and human reports provide a feedback opportunity for organizational-level threat mitigation. See Appendix 2 for the full control pathway in the studied organization.
We review phishing attacks and defenses (§2.1), and then discuss prior empirical studies on phishing training (§2.2).
2.1 Phishing Attacks and Defenses
Phishing attacks pose substantial risks to organizations, who erect both technical and human defenses against them.
2.1.1 Phishing Attacks
In a phishing attack, attackers send messages (phishing lures) to targets whom they wish to persuade to act against their own interests, and the interests of their organizations. Attackers seek results such as sensitive information, account credentials, or the ability to control the victim’s machine [4, 34]. Traditionally, phishing attacks are divided into two classes [25]:
Mass Phishing Campaigns: These use malicious emails sent to thousands or millions of recipients, with common themes such as banking, shipping, and software updates [4]. Individual success rates are low (1-3%) [3], but the massive scale makes them profitable.
Spear Phishing and Targeted Attacks: Spear phishing targets specific individuals or organizations with personalized content from open-source intelligence (OSINT) and reconnaissance. Attackers mine sources such as social media to craft messages that reference real relationships and events [13]. These tailored campaigns have higher success rates (10–20%) [14]. Successful spear phishing attacks include the 2016 Democratic National Committee breach and widespread business email compromise (BEC) schemes [4].
Modern Phishing: Contemporary phishing operations leverage infrastructure including compromised (“trusted”) websites and domain generation algorithms to maintain persistence and evade takedown efforts [3]. In addition, Generative AI is enabling both mass phishing and spear phishing through targeted and fluent emails and other media [4, 80]. LLM-generated phishing content now achieves effectiveness comparable to human experts in organizational settings [9].
@@ 2.1.2 Automated Defenses
Figure˜1 illustrates how these attacks flow through an organization’s email system. Many technical defenses (blue boxes) combat attackers’ behavior, such as SPF, DKIM, and DMARC to verify that mail comes unaltered from authorized infrastructure [51, 22, 52]; ARC for preserving authentication results across forwarding [5]; BIMI for verified logos [10, 12]; and secure email gateways (SEG) for content and header inspection and URL rewriting for time-of-click inspection [1].
After delivery, organizations employ continuous monitoring and feedback loops to adapt controls to evolving campaigns. Organizations blend threat-intelligence updates with AI/ML-driven anomaly scoring and impersonation detection to surface novel lures [58]. For example, security operations center (SOC) analysts triage user-reported messages, promote high-confidence indicators back into blocklists, and tune SEG policies based on observed false positives/negatives. Because adversaries adapt, detection is a dynamic process [4, 3].
See Appendix 2 for the details of the studied organization’s email controls and how user phishing reports create feedback loops that strengthen automated defenses.
2.1.3 Phishing Training as a Defense
@@
Because phishing attacks compromise these technical defenses [14, 13], the human recipients play an important role in phishing defense [30, 15, 2, 28]. There are many approaches to training humans about phishing attacks. Methods range from lecture-style training involving passive videos and readings [56] to simulation feedback providing immediate feedback after interaction with phishing simulations [19, 18]. Interactive training uses gamification or scenario-based learning to engage participants [11, 46], while just-in-time training alerts users in real time during risky behavior [36].
@@
Anti-phishing trainings are often included in the cybersecurity awareness trainings ubiquitous in organizations. Legal mandates drive this practice, e.g., HIPAA requiring security awareness training for healthcare organizations [74]. GDPR and PCI DSS encourage or mandate similar training [76, 21]. Industry standards and frameworks further reinforce these requirements [75], e.g., ISO 27001 calling for human resource security controls including awareness programs [35].
2.2 Empirical Studies of Phishing Training
Table 1: Summary of Prior Experiments on Phishing and Training Interventions. Studies are organized by research context (Lab vs. Real-World). Context: “Lab” denotes research simulations, where subjects are not in their work roles. “Real-World” denotes studies where participants were employees acting in their job capacity. Complexity measure: Whether the study controlled for the difficulty of the phishing tasks. indicates the use of an open control such as NIST Phish Scale, and indicates no control. Hypotheses: How these works informed the hypotheses in the present study (cf. Table˜4).
[82] (2024) Lab University students 96 H3
[16] (2024) Lab University students 117 (Phish Scale) H4
[18] (2014) Real-world Employees 1,359 H2, H4
[19] (2017) Real-world University students 150 H2, H3
[90] (2022) Real-world Employees 119 H2
[53] (2022) Real-world Employees 19,000+ H2
[47] (2023) Real-world Finance employees 5,000 H2
[54] (2024) Real-world Employees 4,554 H2, H3
[29] (2024) Real-world Banking employees “Thousands” H2
[48] (2025) Real-world Healthcare employees 19,500+ H2
Ours Real-world Fintech org. 12,511 (Phish Scale)
Our review and analysis of prior work is summarized in Table˜1. We contribute to the body of knowledge on phishing training by conducting another large-scale study situated in a real-world context (table columns 2-4), incorporating a complexity measure (column 5), and a comparison of multiple training approaches (column 6). Our study tests hypotheses derived from prior work (final column). The remainder of this section describes these contributions. For general surveys of knowledge about human factors in phishing training, we refer the reader to Marshall et al. [56] and Franz et al. [37].
Lab vs. Real-world Studies: Many early phishing training studies were conducted in controlled laboratory settings, where researchers benefit from a high degree of experimental control. Studies of this sort have provided numerous insights, including that active learning and immediate feedback can improve short-term detection rates, and have explored how individual differences affect training outcomes [90, 56]. However, as noted by Franz et al. [37], these controlled settings may lack ecological validity. In laboratory environments, participant samples are typically small and consist of homogeneous groups such as academics and students.
In response, a growing number of studies have begun examining phishing training in real-world organizational environments. For example, Doing et al. [29] analyzed phishing behaviors in bank employees, and Ho et al. [48] studied healthcare employees. Lain et al. has also conducted large-scale real-world measurements of training effectiveness [53, 54] Our study echoes some of these prior results in a different organizational context.
Controlling for Lure Complexity: These studies consistently report little effectiveness of anti-phishing training. These works have not, however, incorporated a standardized method for assessing the difficulty of the phishing lures they use [48, 29]. We contribute the first real-world measurement of the effect of lure complexity on subjects’ interaction behaviors, based in NIST’s Phish Scale.
To assess phishing lure complexity, the NIST Phish Scale categorizes them along two dimensions (Table˜2). Phishing Cues represent observable errors or inconsistencies in the email that might alert users to its fraudulent nature, such as spelling mistakes, suspicious URLs, or formatting irregularities [16]. Premise Alignment measures the relevance and familiarity of the email’s subject matter to the recipient’s typical email environment and organizational context [6]. Higher alignment messages with fewer cues are the most deceptive [14]. The premise of Phish Scale is to allow organizations to benchmark lure complexity and contextualize simulation outcomes [17]. Barrientos et al. [8] conducted early validation work on the NIST Phish Scale, while Dawkins and Jacobs [27] provided the official implementation guidelines. Ours is the first study to validate the Phish Scale’s premise at scale.
Training Approaches: Franz et al. [37] taxonomize five categories of phishing interventions: training, awareness, education, notifications, and design interventions. Most studies have focused on training, which can be divided into lecture-style annual training and embedded training (e.g., through ongoing phishing simulations sent by the organization’s security team). Like some prior studies, our study performs a controlled experiment to compare the effectiveness of two forms of phishing training.
Table 2: NIST Phish Scale for Detection Difficulty [87, 27].